# Reconciling Reality through Simulation: A Real-to-Sim-to-Real Approach for Robust Manipulation

### v2

@@ Abstract
texs/intro2
I Related Work
@@
II RialTo: A Real-to-Sim-to-Real System for Robust Robotic Manipulation
Figure 1: Overview of the real-to-sim pipeline for transfering scenes to the simulator. The first stage consists of scanning the environment, using off-the-shelf tools such as NeRFStudio, ARCode, or Polycam. Each has its strengths and weaknesses and should be used appropriately (see Appendix XI for recommendations). The second stage consists of uploading the reconstructed scene into RialTo’s GUI where the user can cut the mesh, specify joints, and organize the scene as desired. Once complete, the scene can be downloaded as a USD asset, which can be directly imported into the simulator.
II-A System Overview
Our goal is to obtain a control policy that maps real-world sensory observations to robot actions. We only assume access to a small set of demonstrations $(\sim 15)$ containing (observation, action) trajectories collected by an expert, although in principle RialTo can also be used to robustify large, expressive pretrained models as well. Our approach robustifies real-world imitation learning policies using simulation-based RL to make learned controllers robust to disturbances and distractors not present in the demos. The proposed pipeline, RialTo, achieves this with four main steps (Fig LABEL:fig:teaser):
@@ II-A System Overview
The following sections describe each component in detail, along with a full system overview in Fig LABEL:fig:teaser.
@@
II-B Real-to-Sim Transfer for Scalable Scene Generation
@@
While there are various automated techniques for automatically segmenting and adding articulations to meshes [28], in this work, we take a simple human-centric approach. We offer a simple graphical interface for humans to quickly separate meshes and add articulations (see Fig. 1). Our GUI allows users to upload their own meshes and drag/drop, reposition, and reorient them in the global scene. Users can then separate meshes and add joints between different mesh elements, allowing objects like drawers, fridges, and cabinets to be scanned and processed. Importantly, our interface is lightweight, intuitive, and requires minimal domain-specific knowledge. We conducted a study (Section V) evaluating six non-expert users’ experiences with the GUI and found they could scan complex scenes and populate them with a couple of articulated objects in under 15 minutes of active interaction time. Examples of real-world environments with their corresponding digital twins are shown in Fig 3 and Appendix Fig. 15.
The next question is —how do we infer the physics parameters that faithfully replicate the real world? While accurately identifying physical parameters is possible, this can be challenging without considerable interaction [5, 65]. While adapting to dynamics variations is an important direction for future work, in this system we set a single default value for mass and friction uniformly across objects and compensate for the sim-to-real gap to actual real-world values by constraining the learned policy to be close to a small number of real-world demonstrations as discussed in Section II-C.
@@
II-C Robustifying Real-World Imitation Learning Policies in Simulation
Given the simulation environment generated in Section II-B, the next step in RialTo involves learning a robust policy in simulation that can solve desired tasks from a wide variety of configurations and environmental conditions. While this can be done by training policies from scratch in simulation, this is often a prohibitively slow process, requiring considerable manual engineering. Instead, we will adopt a fine-tuning-based approach, using reinforcement learning in simulation to fine-tune a policy initialized from a small number of expert demonstrations collected in the real world. Since training RL directly from visual observations is challenging, we would ideally like to finetune simulation policies that are based on a privileged Lagrangian state. However, real-world demonstrations do not have access to the low-level state information in the environment. To enable the bootstrapping of RL finetuning in simulation from a privileged state using real-world demonstrations, we introduce a novel “inverse distillation” (Section II-C1) procedure that is able to take real-world demonstrations with only raw sensor observations and actions and transfer them to simulation demonstrations, complete with low-level privileged state information. These privileged information demonstrations can then be used to instantiate an efficient RL-based fine-tuning procedure (Section II-C2) in simulation to massively improve policy robustness.
II-C1 Inverse-distillation from Real-to-Sim for Privileged Policy Transfer
We assume a human provides a small number of demonstrations in the real world $\mathcal{D}_{\text{real}}=\{(o_{1}^{i},a_{1}^{i}),\dots,(o_{H}^{i},a_{H}^{i})\}_{i=1}^{N}$, where trajectories contain observations $o$ (3D point clouds) and actions $a$ (delta end-effector pose). Considering that simulation-based RL fine-tuning is far more efficient and performant when operating from a compact state representation [30, 11] (see Section IV-C) and we wish to use real-world human demonstrations to avoid the difficulties with running RL from scratch (see Section IV-B), we want to transfer our observation-action demonstrations from the real world to simulation in a way that allows for subsequent RL fine-tuning in simulation from compact state-based representations. This presents a challenge because we do not have an explicit state estimation system that provides a Lagrangian state for the collected demonstrations in the real world. We instead introduce a procedure, called “inverse-distillation”, for converting our real-world set of demonstrations into a set of trajectories in simulation that are paired with privileged low-level state information.
@@ II-C1 Inverse-distillation from Real-to-Sim for Privileged Policy Transfer
II-C2 Reinforcement Learning Fine-tuning in Simulation
Given the privileged information dataset $\mathcal{D}_{\text{sim}}$, and the constructed simulation environment the goal is to learn a robust policy $\pi^{*}_{\text{sim}}(a|s)$ using reinforcement learning. There are two key challenges in doing so in a scalable way: (1) resolving exploration challenges with minimal reward engineering, and (2) ensuring the policy learns behaviors that will transfer to the real world. ‘ We find that both challenges can be addressed by a simple demonstration augmented reinforcement learning procedure [56, 44, 52], using the Lagrangian state-based dataset $\mathcal{D}_{\text{sim}}$. To avoid reward engineering, we define a simple reward function that detects if the scene is in a desired goal state (detailed sparse reward functions used in each task in Appendix VII). We build on the proximal policy optimization [55] algorithm with the addition of an imitation learning loss as follows (where $\hat{A}_{t}$ is the estimator of the advantage function at step $t$ [55], and $V_{\phi}$ is the learned value function):
@@ II-C2 Reinforcement Learning Fine-tuning in Simulation
In addition to mitigating issues associated with exploration [44, 52], leveraging the additional imitation learning term in the objective helps bias the policy toward physically plausible, safe solutions that improve transfer of behaviors to reality. During this process, we can train the policy for robustness by randomizing initial robot/object/goal poses. Appendix VII contains complete details of our training procedure. The result is a robust policy $\pi^{*}_{\text{sim}}(a|s)$ operating from Lagrangian state that is successful from a wide variety of configurations and environmental conditions.
II-D Teacher-Student Distillation with Co-Training on Real-World Data for Sim-to-Real Transfer
@@
Here the first term corresponds to DAgger training in simulation, while the second term co-trains on real-world expert data. This allows the policy to take advantage of small amounts of high-quality real-world data to bridge the perceptual gap between simulation and real-world scenes and improve generalization compared to only using the data from simulation. We empirically demonstrate (Section II-D) that this significantly increases the resulting success rate in the real world. On a practical note, we refer the reader to Appendix VIII for additional details on the student-teacher training scheme that enables it to be successful in the proposed problem setting.
@@ III Experimental Evaluation
To answer these questions, we evaluate RialTo in eight different tasks, shown in Figure 3 and 7. These include 6-DoF grasping and reorientation of free objects (book on a shelf, plate on a rack, mug on a shelf) and 6-DoF grasping and interacting with articulated objects (drawer and cabinet) on a tabletop and opening a toaster, plate on a rack, putting a cup in the trash in more uncontrolled scenes. More details on the tasks such as their sparse reward functions and randomization setups are presented in Appendix VII. For each task, we consider three different disturbance levels in increasing order of difficulty (see Appendix VII for more details):
@@
We conduct our experiments on a Franka Panda arm with the default parallel jaw gripper, using 6 DoF Cartesian end effector position control. For perceptual inputs, we obtain 3D point cloud observations from a single calibrated depth camera. More details on the hardware setup can be found in Appendix IX. All of the results in the real world are evaluated using the best policy obtained for each method, we report the average across at least 10 rollouts and the bootstrapped standard deviation. Videos of highlights and evaluation runs are available in the website.
Throughout the next sections, we will evaluate RialTo against the following set of baselines and ablations: 1) Imitation learning from 15 and 50 demos (Section III-A); 2) No co-training on real-world data (Section III-B); 3) Co-training on demonstrations in simulation (Section III-B); 4) RialTo from simulation demos (Section III-C2); 5) Learning from an untargeted set of simulated assets (Section III-C1); 6) RialTo without distractors (Section IV-A); 7) RialTo without demos (Section IV-B)
III-A RialTo Learns Robust Policies via Real-to-Sim-to-Real
@@
In this section, we aim to understand whether RialTo can solve complex tasks, showing robustness to variations in configurations, disturbances, and distractors. We compare our approach of real-to-sim-to-real RL fine-tuning against a policy trained only on real-world demos via standard imitation learning (BC). We report the results of running RialTo’s pipeline starting from 15 demos collected directly in simulation and co-training with 15 real-world demos during the teacher-student distillation. In Section III-C2 we show a comparison of running RialTo uniquely on real or sim demos.
@@
Figure LABEL:fig:teaser, 9 and the videos in the website qualitatively show how the resulting policies are robust to many kinds of environment perturbations, including moving the robot, moving the manipulated object and target positions, and adding visual distractors that cause occlusion and distribution shift. The policy rollouts also demonstrate error recovery capabilities, correcting the robot’s behavior in closed loop when, e.g., objects are misaligned or a grasp must be reattempted. This highlights that RialTo provides robustness that does not emerge by purely learning from demonstrations.
@@
We also compare RialTo against behavior cloning with 50 demonstrations to show that the problem is not simply one of slightly more data. Collecting the 50 demonstrations takes a total time of 1 hour and 45 minutes, which is significantly more than the human effort for RialTo for which we collect 15 demos, in 30 minutes, and build the environment in 15 minutes of active time (see Section V). Although more data improves the performance of direct imitation learning from 10% to 40%, 0% to 30%, and 0% to 20% for the three different levels of robustness, the results in Table I show that RialTo achieves approximately 2.5 times higher success rate than pure BC, despite using less than one third the number of demonstrations and taking less than half of the human supervision’s time.
III-B Impact of Co-Training with Real-World Data

### v3

@@ Abstract
Figure 1: RialTo system overview. 1) Transfer the real-world scene to the simulator through an easy-to-use API (see Section III-B). 2) Transfer a policy learned from real-world demonstrations to collect a set of demonstrations with privileged information in simulation. We note this step is optional, and RialTo is compatible with skipping this step and providing demonstrations in simulation (see Section IV-C2) 3) Use the collected set of demonstrations to bias exploration in the RL fine-tuning with sparse rewards of a state-based policy (see Section III-C) 4) Perform teacher-student distillation and deploy the policy in the real world obtaining robust behaviors (see Section III-D).
I Introduction
Imagine a robot that can de-clutter kitchens by putting dishes on a dish rack. Consider all the environmental variations that might be encountered: different configurations of plates or changes in rack positions, a plate unexpectedly slipping in the gripper during transit, and visual distractions, including clutter and lighting changes. For the robot to be effective, it must robustly solve the task across the various scene and object perturbations, without being brittle to transient scene disturbances. Our desiderata is a framework that makes it easy for humans to program the robot to achieve a task robustly under these variations or disturbances. To be a scalable choice for deployment, the framework should not make task-specific assumptions and must seamlessly apply to many tasks.
To design these types of robust robot controllers, one could attempt to train policies across a massive range of scenes and with highly variable objects [12, 21]. This is hard-pressed to provide a scalable solution to robotic learning for two reasons - (1) it is challenging to actually collect or synthesize data across a massive range of scenes since content creation can be challenging in simulation and data collection can be challenging for the real world, (2) a widely general, robust policy may be overly conservative, lowering its performance on the specific target domains encountered on deployment. Alternatively, we suggest that to maximally benefit a specific user, it is more critical that the robot achieves high success in their particular home environment, showing robustness to various local disturbances and distractors that might be encountered in this setting. With this in mind, our goal is to develop a robot learning technique that requires minimal human effort to synthesize visuomotor manipulation controllers that are extremely robust for task performance in deployment environments. The question becomes - how do we acquire these robust controllers without requiring prohibitive amounts of effort for data collection or simulation engineering?
A potential technique for data-driven learning of robotic control policies is to adopt the paradigm of imitation learning (IL), learning from expert demonstration data [58, 56, 25]. However, controllers learned via imitation learning tend to exhibit limited robustness unless a large number of demonstrations are collected. Furthermore, imitation learning does not learn to recover from mistakes or out-of-distribution disturbances unless such behaviors were intentionally demonstrated. This makes direct imitation learning algorithms unsuitable for widespread, robust deployment in real-world scenarios.
The alternative paradigm of reinforcement learning (RL) allows robots to train on self-collected data, reducing the burden on humans for extensive data collection [31] and to discover robust recovery behaviors beyond a set of pre-collected demonstrations (e.g., re-grasping when an object is dropped, re-aligning when an object moves in the gripper, adjusting to external perturbations, etc. — see examples in Fig. 1). However, directly performing RL in the real world is prohibitively slow, often results in unsafe data collection, and is challenging due to problems like resets and reward specification [75]. Therefore, currently, it’s impractical in many cases to employ RL for learning robust control policies directly in the real world. Simulation, on the other hand, offers the ability to collect significant amounts of data broadly, cheaply, safely, and with privileged information  [12, 34, 61, 40, 1]. However, manually constructing geometrically, visually, and physically realistic simulation environments for problems like robotic manipulation in the home can be time and labor-intensive, making it an impractical alternative at scale.
To safely and efficiently learn robust manipulation behaviors, our key insight is to train RL controllers on quickly constructed simulation scenes. By leveraging a video from the target deployment domain, we can obtain scenes complete with accurate geometry and articulation that reflect the appearance and kinematics of the real world. These “in-domain” simulation environments can serve as a sandbox to safely and quickly learn robust policies across various disturbances and distractors, without requiring expensive exploration in the real world. We show how imitation learning policies trained with small numbers of real-world demonstrations can be robustified via large-scale RL fine-tuning in simulation on these constructed simulation environments, using minimal amounts of human effort in terms of environment design and reward engineering. To remove the burden of reward engineering, we leverage a set of real-world demonstrations that bootstrap efficient fine-tuning with reinforcement learning. These real-world demonstrations help narrow the sim-to-real gap and increase the performance of our policies, as shown in Section IV-B. However, transferring real-world demonstrations into simulation is non-trivial because we do not have access to the Lagrangian state of the environment (e.g. object poses). We therefore propose a new “inverse-distillation” technique that enables transferring real-world demos into the simulation. After using RL in constructed simulation environments to robustify the real-world imitation learning policies, the fine-tuned policies can be transferred back to the real world with significantly improved success rates and robustness to test-time disturbances.
Overall, our pipeline simultaneously improves the effectiveness of both reinforcement and imitation learning. RL in simulation helps make imitation learning policies deployment-ready without requiring prohibitive amounts of unsafe, interactive data collection in the real world. At the same time, bootstrapping from real-world demonstration data via inverse distillation makes the exploration problem tractable for RL fine-tuning in simulation. This minimizes the amount of task-specific engineering required by algorithm designers such as designing the dense rewards or manually designing the scenes.
Concretely, we propose RialTo, a system for robustifying real-world imitation learning policies without requiring significant human effort, by constructing realistic simulation analogs for real-world environments on the fly and using these for robust policy learning. Our contributions include:
  • A simple policy learning pipeline that synthesizes controllers to perform diverse manipulation tasks in the real world that (i) reduces human effort in constructing environments and specifying rewards, (ii) produces robust policies that transfer to real-world, cluttered scenes, showing robustness to disturbances and distractors, (iii) requires minimal amounts of expensive and unsafe data collection in the real world.
  • A novel algorithm for transferring demonstrations from the real world to the reconstructed simulation to bootstrap efficient reinforcement learning from the low-level Lagrangian state for policy fine-tuning. We show that this real-to-sim transfer of human demonstrations both improves efficiency and biases policies toward realistic behavior in simulation which effectively transfers back to the real world.
  • An intuitive graphical interface for quickly scanning and constructing digital twins of real-world scenes with articulation, separated objects, and accurate geometries.
[…]
II Related Work
@@
III RialTo: A Real-to-Sim-to-Real System for Robust Robotic Manipulation
Figure 2: Overview of the real-to-sim pipeline for transfering scenes to the simulator. The first stage consists of scanning the environment, using off-the-shelf tools such as NeRFStudio, ARCode, or Polycam. Each has its strengths and weaknesses and should be used appropriately (see Appendix XII for recommendations). The second stage consists of uploading the reconstructed scene into RialTo’s GUI where the user can cut the mesh, specify joints, and organize the scene as desired. Once complete, the scene can be downloaded as a USD asset, which can be directly imported into the simulator.
III-A System Overview
Our goal is to obtain a control policy that maps real-world sensory observations to robot actions. We only assume access to a small set of demonstrations $(\sim 15)$ containing (observation, action) trajectories collected by an expert, although in principle RialTo can also be used to robustify large, expressive pretrained models as well. Our approach robustifies real-world imitation learning policies using simulation-based RL to make learned controllers robust to disturbances and distractors not present in the demos. The proposed pipeline, RialTo, achieves this with four main steps (Fig 1):
@@ III-A System Overview
The following sections describe each component in detail, along with a full system overview in Fig 1.
@@
III-B Real-to-Sim Transfer for Scalable Scene Generation
@@
While there are various automated techniques for automatically segmenting and adding articulations to meshes [29], in this work, we take a simple human-centric approach. We offer a simple graphical interface for humans to quickly separate meshes and add articulations (see Fig. 2). Our GUI allows users to upload their own meshes and drag/drop, reposition, and reorient them in the global scene. Users can then separate meshes and add joints between different mesh elements, allowing objects like drawers, fridges, and cabinets to be scanned and processed. Importantly, our interface is lightweight, intuitive, and requires minimal domain-specific knowledge. We conducted a study (Section VI) evaluating six non-expert users’ experiences with the GUI and found they could scan complex scenes and populate them with a couple of articulated objects in under 15 minutes of active interaction time. Examples of real-world environments with their corresponding digital twins are shown in Fig 4 and Appendix Fig. 16.
The next question is —how do we infer the physics parameters that faithfully replicate the real world? While accurately identifying physical parameters is possible, this can be challenging without considerable interaction [5, 69]. While adapting to dynamics variations is an important direction for future work, in this system we set a single default value for mass and friction uniformly across objects and compensate for the sim-to-real gap to actual real-world values by constraining the learned policy to be close to a small number of real-world demonstrations as discussed in Section III-C.
@@
III-C Robustifying Real-World Imitation Learning Policies in Simulation
Given the simulation environment generated in Section III-B, the next step in RialTo involves learning a robust policy in simulation that can solve desired tasks from a wide variety of configurations and environmental conditions. While this can be done by training policies from scratch in simulation, this is often a prohibitively slow process, requiring considerable manual engineering. Instead, we will adopt a fine-tuning-based approach, using reinforcement learning in simulation to fine-tune a policy initialized from a small number of expert demonstrations collected in the real world. Since training RL directly from visual observations is challenging, we would ideally like to finetune simulation policies that are based on a privileged Lagrangian state. However, real-world demonstrations do not have access to the low-level state information in the environment. To enable the bootstrapping of RL finetuning in simulation from a privileged state using real-world demonstrations, we introduce a novel “inverse distillation” (Section III-C1) procedure that is able to take real-world demonstrations with only raw sensor observations and actions and transfer them to simulation demonstrations, complete with low-level privileged state information. These privileged information demonstrations can then be used to instantiate an efficient RL-based fine-tuning procedure (Section III-C2) in simulation to massively improve policy robustness.
III-C1 Inverse-distillation from Real-to-Sim for Privileged Policy Transfer
We assume a human provides a small number of demonstrations in the real world $\mathcal{D}_{\text{real}}=\{(o_{1}^{i},a_{1}^{i}),\dots,(o_{H}^{i},a_{H}^{i})\}_{i=1}^{N}$, where trajectories contain observations $o$ (3D point clouds) and actions $a$ (delta end-effector pose). Considering that simulation-based RL fine-tuning is far more efficient and performant when operating from a compact state representation [32, 11] (see Section V-C) and we wish to use real-world human demonstrations to avoid the difficulties with running RL from scratch (see Section V-B), we want to transfer our observation-action demonstrations from the real world to simulation in a way that allows for subsequent RL fine-tuning in simulation from compact state-based representations. This presents a challenge because we do not have an explicit state estimation system that provides a Lagrangian state for the collected demonstrations in the real world. We instead introduce a procedure, called “inverse-distillation”, for converting our real-world set of demonstrations into a set of trajectories in simulation that are paired with privileged low-level state information.
@@ III-C1 Inverse-distillation from Real-to-Sim for Privileged Policy Transfer
III-C2 Reinforcement Learning Fine-tuning in Simulation
Given the privileged information dataset $\mathcal{D}_{\text{sim}}$, and the constructed simulation environment the goal is to learn a robust policy $\pi^{*}_{\text{sim}}(a|s)$ using reinforcement learning. There are two key challenges in doing so in a scalable way: (1) resolving exploration challenges with minimal reward engineering, and (2) ensuring the policy learns behaviors that will transfer to the real world. ‘ We find that both challenges can be addressed by a simple demonstration augmented reinforcement learning procedure [60, 46, 54], using the Lagrangian state-based dataset $\mathcal{D}_{\text{sim}}$. To avoid reward engineering, we define a simple reward function that detects if the scene is in a desired goal state (detailed sparse reward functions used in each task in Appendix VIII). We build on the proximal policy optimization [59] algorithm with the addition of an imitation learning loss as follows (where $\hat{A}_{t}$ is the estimator of the advantage function at step $t$ [59], and $V_{\phi}$ is the learned value function):
@@ III-C2 Reinforcement Learning Fine-tuning in Simulation
In addition to mitigating issues associated with exploration [46, 54], leveraging the additional imitation learning term in the objective helps bias the policy toward physically plausible, safe solutions that improve transfer of behaviors to reality. During this process, we can train the policy for robustness by randomizing initial robot/object/goal poses. Appendix VIII contains complete details of our training procedure. The result is a robust policy $\pi^{*}_{\text{sim}}(a|s)$ operating from Lagrangian state that is successful from a wide variety of configurations and environmental conditions.
III-D Teacher-Student Distillation with Co-Training on Real-World Data for Sim-to-Real Transfer
@@
Here the first term corresponds to DAgger training in simulation, while the second term co-trains on real-world expert data. This allows the policy to take advantage of small amounts of high-quality real-world data to bridge the perceptual gap between simulation and real-world scenes and improve generalization compared to only using the data from simulation. We empirically demonstrate (Section III-D) that this significantly increases the resulting success rate in the real world. On a practical note, we refer the reader to Appendix IX for additional details on the student-teacher training scheme that enables it to be successful in the proposed problem setting.
@@ IV Experimental Evaluation
To answer these questions, we evaluate RialTo in eight different tasks, shown in Figure 4 and 8. These include 6-DoF grasping and reorientation of free objects (book on a shelf, plate on a rack, mug on a shelf) and 6-DoF grasping and interacting with articulated objects (drawer and cabinet) on a tabletop and opening a toaster, plate on a rack, putting a cup in the trash in more uncontrolled scenes. More details on the tasks such as their sparse reward functions and randomization setups are presented in Appendix VIII. For each task, we consider three different disturbance levels in increasing order of difficulty (see Appendix VIII for more details):
@@
We conduct our experiments on a Franka Panda arm with the default parallel jaw gripper, using 6 DoF Cartesian end effector position control. For perceptual inputs, we obtain 3D point cloud observations from a single calibrated depth camera. More details on the hardware setup can be found in Appendix X. All of the results in the real world are evaluated using the best policy obtained for each method, we report the average across at least 10 rollouts and the bootstrapped standard deviation. Videos of highlights and evaluation runs are available in the website.
Throughout the next sections, we will evaluate RialTo against the following set of baselines and ablations: 1) Imitation learning from 15 and 50 demos (Section IV-A); 2) No co-training on real-world data (Section IV-B); 3) Co-training on demonstrations in simulation (Section IV-B); 4) RialTo from simulation demos (Section IV-C2); 5) Learning from an untargeted set of simulated assets (Section IV-C1); 6) RialTo without distractors (Section V-A); 7) RialTo without demos (Section V-B)
IV-A RialTo Learns Robust Policies via Real-to-Sim-to-Real
@@
In this section, we aim to understand whether RialTo can solve complex tasks, showing robustness to variations in configurations, disturbances, and distractors. We compare our approach of real-to-sim-to-real RL fine-tuning against a policy trained only on real-world demos via standard imitation learning (BC). We report the results of running RialTo’s pipeline starting from 15 demos collected directly in simulation and co-training with 15 real-world demos during the teacher-student distillation. In Section IV-C2 we show a comparison of running RialTo uniquely on real or sim demos.
@@
Figure 1, 10 and the videos in the website qualitatively show how the resulting policies are robust to many kinds of environment perturbations, including moving the robot, moving the manipulated object and target positions, and adding visual distractors that cause occlusion and distribution shift. The policy rollouts also demonstrate error recovery capabilities, correcting the robot’s behavior in closed loop when, e.g., objects are misaligned or a grasp must be reattempted. This highlights that RialTo provides robustness that does not emerge by purely learning from demonstrations.
@@
We also compare RialTo against behavior cloning with 50 demonstrations to show that the problem is not simply one of slightly more data. Collecting the 50 demonstrations takes a total time of 1 hour and 45 minutes, which is significantly more than the human effort for RialTo for which we collect 15 demos, in 30 minutes, and build the environment in 15 minutes of active time (see Section VI). Although more data improves the performance of direct imitation learning from 10% to 40%, 0% to 30%, and 0% to 20% for the three different levels of robustness, the results in Table I show that RialTo achieves approximately 2.5 times higher success rate than pure BC, despite using less than one third the number of demonstrations and taking less than half of the human supervision’s time.
IV-B Impact of Co-Training with Real-World Data